OpenStudy (anonymous):

x 11 1 2 −3 f(x) −1209 1 −3 37 g(x) 2432 2 11 −74 f′(x) −341 −1 −8 −33 g′(x) 683 3 17 67 (fg)′(2)=? f(2)/(g(2)+5)=? (fg)(1)=? (f−g)′(2)=?

OpenStudy (anonymous):

x 11 1 2 −3 f(x) −1209 1 −3 37 g(x) 2432 2 11 −74 f′(x) −341 −1 −8 −33 g′(x) 683 3 17 67 (fg)′(2)=? f(2)/(g(2)+5)=? (fg)(1)=? (f−g)′(2)=?